OpenStudy (anonymous):

What is the midpoint for the coordinates (-2, 4) and (6, 12)? (2, 8) (4, 8) (8, 2) (2, 4)

OpenStudy (help!!!!):

use the midpoint formula

OpenStudy (anonymous):

midpoint?? formula??

OpenStudy (anonymous):

2,8

OpenStudy (help!!!!):

yeah just google it seriously

OpenStudy (help!!!!):

Like its better to get information from a valid source

OpenStudy (agreene):

\[\left( \frac{x_2+x_1}2,\frac{y_2+y_1}2 \right)\] is the midpoint formula